OpenStudy (anonymous):

Two point charges are placed on the x-axis as follows: charge q1= 3.96 is located at x=0.202 , and charge q2= 4.99 is at x=-0.302 .What is the magnitude of the total force exerted by these two charges on a negative point charge q3= -5.95 that is placed at the origin?

OpenStudy (egenriether):

The force at the origin will be the sum of each individual one. One force will be given by Kc*(Q1*Q2)/R^2. Where Kc = 1/(4*pi*eps_0) Note the Q's carry the sign. Since these are all on the x axis, the vector components aren't hard but the direction matters. A negative force means they are attracted, as these will be. Get each answer and add them up.

OpenStudy (anonymous):

the total force on the negative charge placed at the origin=F1+F2 where \[F1=-(3.96*5.95)/(4*3.14*8.854e-12*0.202*0.202)\] F2=−(4.99∗5.95)/(4∗3.14∗8.854e−12∗0.302∗0.302) Now F1 and F2 are opposite to each other and we have the three charges in the x-axis so it is straight away to find the difference between the two forces F1 and F2 ie.F1-F2 resultant force is F=5.193-2.928 =2.265*10^12 newton (magnitude) If we consider the given numericals in the SI units, charge in coulombs, distance in metres. The direction of the resultant force will be towards q1 and as a result the charge q3 is attracted towards q1.